OpenStudy (anonymous):

Simplify the expression -8 x 5 + 16 − 8. Select one of the options below as your answer: A. -104 B. -32 C. -24 D. 48

OpenStudy (anonymous):

b

OpenStudy (anonymous):

-40+16 = -24 -24-8 = -32

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

plz click on good answer to close the question